OpenStudy (anonymous):

Solve x2 + 6x + 7 = 0.

OpenStudy (anonymous):

x = 2

OpenStudy (anonymous):

I don't know, I'm probably wrong.

OpenStudy (anonymous):

x^2 - 6x - 7 = 0 (x-7)(x+1) So x = 7; x = -1 Check your work: 7^2 - (6)(7) - 7 = 0 49 - 42 - 7 = 0 0 =0 (-1)^2 - (6)(-1) - 7 1 + 6 - 7 = 0 0 = 0 So, they check.

OpenStudy (anonymous):

@biggdaddymiller That answer is wrong

OpenStudy (anonymous):

x = -2

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

@biggdaddymiller I still think that is wrong :( I'm pretty sure I solved it correctly

OpenStudy (anonymous):

yea iz 0=0 me chacked :)

OpenStudy (anonymous):

@KILL4PHUN So I'm right

OpenStudy (anonymous):

yupp.

OpenStudy (anonymous):

@mdf913 My answer is correct :) Happy to help

OpenStudy (anonymous):

A) x=-1 x= -5 B) 3 (plus/minus) square root of 2 C) -3 (plus/minus) square root of 2 D) -3 + Square root of 2 divided by 2

OpenStudy (anonymous):

These are my answer choices @lauren24680

OpenStudy (anonymous):

\[\times\]

OpenStudy (anonymous):

Okay, I'm pretty sure the answer is A

OpenStudy (anonymous):

Thank you so much!! :)

OpenStudy (anonymous):

You're welcome! :)

OpenStudy (anonymous):

I have more if you want to help more. If you don't mind. ._.

OpenStudy (anonymous):

Alright, I'll try it :)

OpenStudy (anonymous):

Which of the following is a solution of x2 + 5x = -2?

OpenStudy (anonymous):

\[5\pm (\square \root) of 33 divided by 2\]

OpenStudy (anonymous):

B) 5 (plus/minus) square root of 17 divided by 2 C) -5 (plus/minus) square root of 33 divided by 2 D) -5 (plus/minus) the square root of 17 divided by 2

OpenStudy (anonymous):

Hmmm

OpenStudy (anonymous):

Just give me a few minutes, then I'll answer

OpenStudy (anonymous):

Okay @mdf913, I'm pretty sure the answer is C